Find S60 for the sequence 45, 51, 57, 63, ... 11 S60Find the sum of the first 60 terms of the arithmetic sequence. 14, 19, 24, 29, ... What is the sum of the first 60 terms?
Answer #1

Please do vote:-

1) Given Artiknetic sequence 45, 51,57, 63,.. First term aizus Common difference d = 51.45=57-5156 Ost To find sum, Sn = 2 (2